• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 14:32:43 on April15,2008

    Entry number 232230

    This entry is a followup to: 232209

    keyword=D3 L-HRS fixed (I think)

    The main problem with D3 BCM signal on L-HRS was that it had been
    fanned out using a splitter at the patch panel on the floor level
    in Hall A. Now I remember this was for Compton FADC DAQ (so you can
    blame me) and the cable went nowhere since that DAQ was dismantled
    in Feb. Hence small reflections from an unterminated cable.

    Another issue is that there is a -128 mV (negative) offset to the
    signal at the input to the scalers. This originates at the V-to-F
    upstairs in the counting room where the TTL output does not go
    to zero in between pulses (i.e. not a legal TTL signal). The
    TTL/NIM converter next to it translates this and the DC offset
    shows up at the scaler.

    If the DC offset is stable I think there is no double pulsing at
    the scaler. But if there's a small reflection which perhaps changes
    in amplitude with time, that would explain what we've been seeing.
    It has to cross a threshold for the scalers to count it.

    A DC offset appears on some other V-to-F outputs in the BCM system
    as well. This is worrisome. The NIM crate power supplies are ok.

    After this fix, Jose and I did the usual continuity check:
    unplug/plug the V-to-F output, verify signals in scalers.

    Ready to run ...